Question 6 Using stock solutions to make up a solution: from final concentration to volume You need to digest a sample of DNA with a restriction enzyme. You will do this reaction later in the semester. The buffer for this reaction has a final concentration of 40 mM Tris, pH 7.5, 10 mM MgCl2 and 50 mM KCl. You have the following stock solutions: 0.5 M Tris, pH 7.5, 1 M MgCl2 and 2 M KCl. How would you make up 100 mL of this buffer?
Answer #1

Considering:

C1V1 = C2V2

1. Tris:

C2 = 40mM

V2 = 100 ml

C1= 0.5 M = 500mM

V1 = Unknown

V1= (100X 40)/ 500 = 8 ml (from stock).

2. MgCl2:

C2 = 10mM

V2 = 100 ml

C1= 1 M = 1000mM

V1 = Unknown

V1= (100X 10)/ 1000 = 1 ml (from stock).

3. KCl:

C2 = 50mM

V2 = 100 ml

C1= 2 M = 2000mM

V1 = Unknown

V1= (100X 50)/ 2000 = 2.5 ml (from stock).

Thus, 100 ml of buffer can be made by taking 8 ml from Tris stock, 1 ml MgCl2 and 2.5 ml KCl

= (8+1+2.5) = 11.5 ml

Making final volume to 100 ml by adding (100 -11.5) ml = 88.5 ml water.